OpenStudy (anonymous):

applying the Commutative Property of Addition to 17.5+24.9+2.5

OpenStudy (anonymous):

the commutative property of addition is a + b = b + a

OpenStudy (anonymous):

so 17.5 + 24.9 + 2.5 = 2.5 + 17.5 + 24.9